OpenStudy (anonymous):

A firm wants to determine the amount of frictional torque in their current line of grindstones, so they can redesign them to be more energy efficient. To do this, they ask you to test the best-selling model, which is basically a diskshaped grindstone of mass 1.6 kg and radius 8.90 cm which operates at 740 rev/min. When the power is shut off, you time the grindstone and find it takes 27.5 s for it to stop rotating.

OpenStudy (anonymous):

What is the frictional torque exerted on the grindstone?

OpenStudy (anonymous):

What I did at first is torque=m*g*R, but I got the wrong answer. Can someone please help me explaining why my way is wrong

OpenStudy (anonymous):

average Torque = change in angular momentum / time time = 27.5 seconds change in angular momentum = 0 - I omega I = moment of inertia of solid disk of mass 1.6 kg and radius .089 meters = (1/2) m r^2 omega = 740 rev/min (1 min/60s)(2 pi radians/rev)